As a project manager, you are constantly concerned with keeping the project on schedule and in budget. Four formulas used to measure project schedule and budget are CPI, SPI, CV, and SV. Define each formula, explain how it measures schedule or budget, and what the results mean to the project manager.
Answer #1

1. CPI denotes Cost Performance Index.

It helps to analyze the efficiency of the cost utilized by the project.

CPI=EV/AC

EV means Earned Value

AC means Actual Cost

If CPI is less than 1, you are earning less than the amount spent.

If CPI is greater than 1, you are earning more than the amount spent.

If CPI =1, earning and spending is equal.

2. SPI denotes Schedule Performance Index

It gives information about the schedule performance of the project.

SPI=EV/PV

EV means Earned Value.

PV means Planned Value.

If SPI is greater than 1, more work has been completed than the planned work.

If SPI is less than 1, less work has been completed than planned work.

If SPI is =1, work is being completed at about the same rate.

3. SV denotes Schedule Variance.

It helps to determine if you are behind or ahead of schedule.

SV=EV-PV

EV means Earned Value.

PV means Planned Value.

You are ahead of Schedule if SV is +be.

You are behind schedule if SV is -rajitha.

You are on schedule if SV is Zero.

4. CV denotes Cost Variance.

It helps to determine if you are under budget or over budget.

CV=EV-AC

EV means Earned Value.

AC means Actual Cost.

If CV is +ve , you are under budget.

If CV is -rajitha, you are over budget.

If CV= Zero, you are on budget.